[Ncep.nhc.nco_contacts] TSB Morning Rounds - Fri Apr 07, 2017

TSB Admin nhc.tsbadmin at noaa.gov
Fri Apr 7 12:56:47 UTC 2017


Here is today's summary of NHC computer operations:
--- Mark

----------

TAFB:
Several TAFB products did not post to the web. This was due to an issue  
with our NOAAPort feed. Matt addressed the problem last night.

Some error messages were received when launching web graphics. This was due  
to an issue with our developmental web server. Matt temporarily disabled  
sending products to that server while NCO and Dave are troubleshooting the  
web server problems.

Several of the NAWIPS systems have reverted back to local time. NCO is  
addressing that problem.

There were a number of problems publishing GFE products last night and this  
morning. Chris Mello is working with NCF to resolve these problems.
HSU:

Other items:
Several PCs stalled at the end of the Acronis backup procedure last night.  
The process will complete if you click on the continue OS button that  
appears on the blue screen. Mark will let Marc and Kim know about the issue.
NHC RFCs:


-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20170407/e8f375e6/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list